Americium-241 is a radioactive isotope that emits alpha particles. The rate constant for this radioactive decay is 1.47 x 10-3 per year.

a. What is the half-life for the decay of americium-241?

b. How long will it take for 45.0% of the americium to decay?
Answer #1

All radioactive decay processes follows first order kinetics.

First order rate constant is given as, k = I lola at initial concentration to (a-x) → Concentration after time t Half-life &K= 1:47X10-best year 2.) tyg - 0.693 tv = 471-43 year 1.47 x 10-3 b.) a = 100% a-x =100-45 55% K= 11 la |:47x10-P = In 100 t